OpenStudy (anonymous):

factor the expression 3cos^2x+8cosx-3

OpenStudy (danjs):

\[3\cos^2(x) + 8\cos(x) -3\] Let : cos(x) = u \[3u^2 +8u - 3\]

OpenStudy (danjs):

can you factor that last expression?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

ik what to do

OpenStudy (danjs):

\[(3u \pm a)(u \pm b)\] need to figure a and b

OpenStudy (anonymous):

ok

OpenStudy (danjs):

the factors of 3, are only 1 and 3 a and b will be 1 and 3, or 3 and 1

OpenStudy (anonymous):

oh and then you substitute the cis back in

OpenStudy (anonymous):

cos

OpenStudy (danjs):

yeah , after you factor in terms of u, sub back in u = cos(x)

OpenStudy (anonymous):

so (3cosx-1)(cosx+3)

OpenStudy (danjs):

\[(3u -1)(u + 3)\] u = cos(x) \[3\cos^2(x) + 8\cos(x) - 3~~=~~(3\cos(x) - 1)*(\cos(x) +3)\]

OpenStudy (danjs):

yep

OpenStudy (anonymous):

thanx and do you think you can help me with one more problem

OpenStudy (danjs):

sure

OpenStudy (anonymous):

determine the value of B in the equation 3+cos2B=-2sin^2B+7B

OpenStudy (danjs):

\[3 + \cos(2B) = -2\sin^2(B) + 7B\]

OpenStudy (danjs):

is that right

OpenStudy (anonymous):

yea and the B is suppose to be beta

OpenStudy (danjs):

+7 times beta at the end righ

OpenStudy (anonymous):

yea

OpenStudy (danjs):

do you remember the identity \[\cos(2\beta ) = 1-2\sin^2(\beta)\]

OpenStudy (anonymous):

yea from the trig identities

OpenStudy (danjs):

\[3+[1-2\sin^2(\beta)] = -2\sin^2(\beta)+7*\beta\]

OpenStudy (danjs):

both sides now have -2sin^2(Beta) and cancels out leaving just 3 + 1 = 7B 4 = 7B B = 4/7

OpenStudy (anonymous):

wait how did the two -2sin^2 beta cancel

OpenStudy (danjs):

\[4-2\sin^2(\beta) = -2\sin^2(\beta)+7\beta \] add\[+2\sin^2(\beta)\]to both sides

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how do ik when i need to use trig functions identities